By substituting the value of x in the following, first compute y and then represent y in Radix and Diminished-Radix complement formats (for respective Radix),

iii (x)10 + (1011)2 = ( y )8

iv (x)10 × (15)8 = ( y )2 

value of x is 49


1
Expert's answer
2021-06-30T06:01:43-0400

Part iii

(1011)2 = 1110


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 011_{10}
\end{array}}" "\underline{\begin{array}{cc}
 X \\
 + Y
\end{array}}"

Z


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 011_{10}
\end{array}}"

50110

Hence discarding the leading 1 the result is 501.


Part iv

"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 120_{10}
\end{array}}" "\underline{\begin{array}{cc}
 X \\
 + Y
\end{array}}"

Z


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 120_{10}
\end{array}}"

61010

Hence the result is 610